52 well-shuffled playing cards are distributed at random among 4 player's.Find the probability that a hand contain

(i) 3 aces

(ii) 2 kings and one queen

(iii) No cards of diamond

(iv) 5 pictured cards .


1
Expert's answer
2022-02-07T14:05:42-0500
"52\/4=13"

There are "\dbinom{52}{13}=\dfrac{52!}{13!(52-13)!}=635,013,559,600" ways to distribute 13 cards to each of four people.

(i)


"P(3\ aces)=\dfrac{\dbinom{4}{3}\dbinom{52-4}{13-3}}{\dbinom{52}{13}}"

"=\dfrac{4(6,540,715,896)}{635,013,559,600}=0.0412"

(ii)


"P(2\ kings\ and\ 1\ queen)=\dfrac{\dbinom{4}{2}\dbinom{4}{1}\dbinom{52-8}{13-3}}{\dbinom{52}{13}}"

"=\dfrac{6(4)(2,481,256,778)}{635,013,559,600}=0.0938"



(iii)


"P(No\ diamonds)=\dfrac{\dbinom{13}{0}\dbinom{52-13}{13-0}}{\dbinom{52}{13}}"

"=\dfrac{1(8,122,425,444)}{635,013,559,600}=0.0128"



(iv)


"P(5\ picture\ cards)=\dfrac{\dbinom{12}{5}\dbinom{52-12}{13-5}}{\dbinom{52}{13}}"

"=\dfrac{792(76,904,685)}{635,013,559,600}=0.0959"